# softMathematical software

### v1

@@ 4. Iterations of the self-adhesion operator
Find out whether ${\mathcal{M}}\in\mathfrak{st}^{\diamond}(N)\setminus\mathfrak{st}^{\diamond\diamond}(N)$ exists in case $|N|=4$.
5. The case of a small number of variables
In this section we present the results of our computations. We also give a few instructive examples obtained as their results. A summary of numerical data about the lattices we consider in case $|N|=4$ is available in Appendix C.
@@ 5.2. Structural semi-graphoids over 4 variables
Note that our results agree with the analysis presented in [31, § V.A], or with the rules A3–A7 from [28]. In particular, the lattice $(\mathfrak{st}(N),\subseteq)$ is implicatively perfect in case $|N|=4$. Moreover, both the semi-graphoidal rules and the rules (E:1)–(E:5) can be generalized to any higher cardinality than $|N|\geq 4$ [6, 5].
@@ 5.3. Self-adhesive semi-graphoids over 4 variables
We also computed the canonical implicational basis ${\mathbb{G}}_{{\mathcal{F}}}$ for ${\mathcal{F}}=\mathfrak{sg}^{\diamond}(N)$ in case $|N|=4$. It decomposes into 9 permutational types: besides two semi-graphoidal types mentioned in Section 5.1 it has 7 additional types and we divided them into two subgroups. The implications in the first subgroup can be interpreted as weaker versions of probabilistic equivalences from [31, § V.A], recalled in Section 5.2, and, therefore, we use the label SE to denote them:
@@
The implications in the second subgroup can be interpreted as weaker versions of probabilistic implications presented in [31, § V.B], for which reason we use the label SI to denote them:
@@
This is a comparison with the implications from [31, § V.B]:
@@
We also confirmed independently the validity of the rules (SE:1)–(SE:2) and (SI:1)–(SI:5) within the frame $\mathfrak{sg}^{\diamond}$ by finding their certificates in the form of semi-graphoidal derivation sequences, as illustrated in Example 5.1. Specifically, we found out that the antecedent set in (SE:1) is not self-adhesive at sets $ij$ and $jk$, while the antecedent sets in (SE:2) and in the rules labeled by SI are not self-adhesive at the sets $ij$, $ik$, and $jk$.
@@ 5.4. Self-adhesive structural semi-graphoids: 4 variables
Put $N:=\{a,b,c,d\}$. The model :=${\mathcal{M}}:=\{\,ab|d,ac|d,ad|b,bc|\emptyset,bc|d\,\}$ belongs to $\mathfrak{st}^{\diamond}(N)$ while :=${\mathcal{M}}^{\star}:=\{\,ab|c,ac|b,ad|c,bc|a,bc|ad\,\}\not\in\mathfrak{st}^{\diamond}(N)$. The latter fact follows from the rule (SI:2) above where $i:=a$, $j:=c$, $k:=b$, and $\ell:=d$.
@@
{open}
Find our whether the relation $\mathfrak{sg}^{\diamond}(N)\cap\mathfrak{st}(N)=\mathfrak{st}^{\diamond}(N)$ hold in general.
@@ 5.5. The case of 5 variables
In general, the coatoms of this lattice correspond to the extreme rays of the cone of standardized supermodular functions, that is, to supermodular functions $m:{\mathcal{P}}(N)\to{\mathbb{R}}$ satisfying $m(K)=0$ if $K\subseteq N$, $|K|\leq 1$, see [33, Corollary 30]. Thus, the coatoms for $|N|=5$ were obtained by computing the extreme rays of that cone [32]. We raised a couple of questions:
@@
Surprising answers were that only 154 of those types belong to $\mathfrak{sg}^{\diamond}(N)$ and these coincide with the types belonging to $\mathfrak{st}^{\diamond}(N)$. This observation seems to support the conjecture indicated in Task 5.4. Consequences of those results are discussed in detail in Section 6.
5.6. Other examples of abstract CI frames
A semi-graphoid is called compositional if it satisfies the reversed implication, and a compositional graphoid is a semi-graphoid satisfying both implications. It is evident that ${\mathcal{M}}$ is a graphoid if and only if ${\mathcal{M}}^{\star}$ is a compositional semi-graphoid.
The above definitions, thus, delimit the frame $\mathfrak{gr}$ of graphoids, the frame $\mathfrak{gr}^{\star}$ of compositional semi-graphoids, and the frame $\mathfrak{cgr}$ of compositional graphoids. By definition, $\mathfrak{cgr}\subseteq\mathfrak{gr},\mathfrak{gr}^{\star}\subseteq\mathfrak{sg}$. Our notation reflects the fact that $\mathfrak{gr}^{\star}$ can be interpreted as the dual frame to $\mathfrak{gr}$ because ${\mathcal{M}}\in\mathfrak{gr}^{\star}(N)\,\Leftrightarrow\,{\mathcal{M}}^{\star}\in\mathfrak{gr}(N)$ for any variable set $N$ and ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$. It is known that $\mathfrak{gr}\setminus\mathfrak{pr}\neq\emptyset$ and $\mathfrak{pr}\setminus\mathfrak{gr}\neq\emptyset$. In case $|N|=4$, precisely $369$ permutational types of graphoids are probabilistically representable.
A result of our computations was the finding that, in case $|N|=4$, the family $\mathfrak{gr}(N)$ contains 6482 elements, which fall down into 421 permutational types. The number of (meet) irreducible models in $\mathfrak{gr}(N)$ is 408 and they decompose into 32 types. The lattice $(\mathfrak{gr}(N),\subseteq)$ has 14 coatoms of 4 types; thus, it is not coatomistic. The family $\mathfrak{gr}^{\diamond}(N)$ in case $|N|=4$ coincides with $\mathfrak{gr}(N)$.
The family of undirected graphical models is then
This defines a CI frame $\mathfrak{ug}$, which is closed under copying and marginalization, but not under intersection. Well-known facts are that $\mathfrak{ug}\subseteq\mathfrak{pr}\subseteq\mathfrak{sg}$ [14]. The $\mathfrak{ug}$-frame has a finite “axiomatization” [26]. We can now give an example that illustrates the claim in the remark n. 2 after Definition 3.3, namely that there is a pair of frames $\mathfrak{f}_{1}\subseteq\mathfrak{f}_{2}\subseteq\mathfrak{sg}$ closed under marginalization and copying and a model ${\mathcal{M}}\in\mathfrak{f}_{1}(N)$ which is self-adhesive relative to $\mathfrak{f}_{2}$ but not relative to $\mathfrak{f}_{1}$.
Example 5.4.
( self-adhesivity is relative to a frame )
Consider frames $\mathfrak{f}_{1}:=\mathfrak{ug}$, $\mathfrak{f}_{2}:=\mathfrak{sg}$ and the variable set $N:=\{a,b,c\}$. The model :=${\mathcal{M}}:=\{\,bc|a\,\}\subseteq{\mathcal{K}}(N)$ belongs to $\mathfrak{ug}(N)$: it is induced by the graph $G$ obtained from the complete graph over $N$ by the removal of the edge between $b$ and $c$. Then ${\mathcal{M}}\in\mathfrak{sg}^{\diamond}(N)$ but ${\mathcal{M}}\not\in\mathfrak{ug}^{\diamond}(N)$. To this end consider the set $L:=bc$ and a bijection $\varphi:N\to M:=\{\bar{a},b,c\}$ with $\varphi(a):=\bar{a}$, $\varphi(b):=b$, and $\varphi(c):=c$ (see Figure 2). Then ${\mathcal{M}}$ is not self-adhesive at $L$ relative to $\mathfrak{ug}$. Indeed, a potential adhesion ${\mathcal{E}}$ of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$ in $\mathfrak{ug}(NM)$ has to satisfy
@@ 6. Application to entropic region delimitation
Well-known facts are that the entropy function $h_{\mbox{\scriptsize\boldmath$\xi$}}:{\mathcal{P}}(N)\to{\mathbb{R}}$ is submodular, non-decreasing with respect to inclusion, and satisfies $h_{\mbox{\scriptsize\boldmath$\xi$}}(\emptyset)=0$. Thus, in terms of Section 2.2.3, it is (the rank function of) a polymatroid over $N$.
A polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ is then called entropic if $h=h_{\mbox{\scriptsize\boldmath$\xi$}}$ for some discrete random vector $\xi$ over $N$. It is called almost entropic if it is a limit of entropic polymatroids (in the Euclidean topology on ${\mathbb{R}}^{{\mathcal{P}}(N)}$). The entropic region, also named the entropy region, is the set of all entropic polymatroids and the almost entropic region is the set of all almost entropic polymatroids; the latter set is known to be a closed cone in ${\mathbb{R}}^{{\mathcal{P}}(N)}$ [35, § 15.1].
@@
Given a bijection $\varphi:N\to M$, the $\varphi$-copy of a polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ is the polymatroid $h\circ\varphi_{-1}\in{\mathbb{R}}^{{\mathcal{P}}(M)}$. A pair of sets $N,M$ is called modular in a polymatroid $\hat{h}\in{\mathbb{R}}^{{\mathcal{P}}(NM)}$ if $\hat{h}(N)+\hat{h}(M)=\hat{h}(NM)+\hat{h}(N\cap M)$. A polymatroid $\hat{h}\in{\mathbb{R}}^{{\mathcal{P}}(NM)}$ extends $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ if $h$ is the restriction of $\hat{h}$ to ${\mathcal{P}}(N)$. We say that a polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ is self-adhesive at a set $L\subseteq N$ if,
@@
such that N,M is a modular pair in $~{}~{}\mbox{such that $N,M$ is a modular pair in $\hat{h}$}\,.$
A polymatroid over $N$ is self-adhesive if it is self-adhesive at any $L\subseteq N$.
Recall from Section 2.2.3 that a polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ induces a CI model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ as follows: $ij|K\in{\mathcal{M}}$ if and only if $\Delta h(ij|K)=0$.
@@
Proof 6.3.
To show (a) is true, consider a set $L\subseteq N$ and the respective bijection $\varphi:N\to M$. Realize that the set of polymatroids $\hat{h}\in{\mathbb{R}}^{{\mathcal{P}}(NM)}$ in which $N,M$ is a modular pair and that satisfy $\hat{h}(A)=\hat{h}(\varphi(A))$ for $A\subseteq N$ is a polyhedral cone. A polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ is self-adhesive at $L$ if and only if it belongs the projection of that set of vectors $-^\hat{h}$ from ${\mathbb{R}}^{{\mathcal{P}}(NM)}$ to ${\mathbb{R}}^{{\mathcal{P}}(N)}$. Basic facts from polyhedral geometry are that the class of polyhedral cones is closed under projections and finite intersections. The application of these facts yields (a).
For (b) we first realize that any entropic polymatroid is self-adhesive. The arguments for this from [23, Corollary 1] are analogous to those in our proof of Lemma 3.7. Then (a) implies that set of self-adhesive polymatroids over $N$ is closed, which forces that it contains the closure of the entropic region.
As concerns (c), recall from the remark n. 1 in the end of Section 2.2.3 that ${\mathcal{M}}\in\mathfrak{st}(N)$ if and only if it is induced by a polymatroid over $N$. Given ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ induced by a self-adhesive polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$, one needs to observe that ${\mathcal{M}}$ is self-adhesive at (any) $L\subseteq N$ relative to $\mathfrak{st}$. Having fixed the corresponding bijection $\varphi:N\to M$, take the polymatroid $\hat{h}\in{\mathbb{R}}^{{\mathcal{P}}(NM)}$ from Definition 6.1 and it is straightforward to verify that it induces ^$\hat{{\mathcal{M}}}\in\mathfrak{st}(NM)$ which is an adhesive extension of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$.
@@
Proof 6.4.
Realize that any facet-defining inequality for the cone of self-adhesive polymatroids is a valid information inequality.
To interpret the results from Section 5.5 we recall some geometric facts. A polymatroid $h$ over $N$, $|N|\geq 1$, is called tight if $h(N)=h(N\setminus i)$ for any $i\in N$ and modular if $\Delta h(ij|K)=0$ for any $ij|K\in{\mathcal{K}}(N)$. Every polymatroid $h$ then decomposes into its modular part and its tight part $\tilde{h}$ inducing the same CI model as $h$. Moreover, there exists a linear one-to-one mapping between tight polymatroids and standardized supermodular functions which preserves the induced CI model, see [33, formulas (32) in § 7.2]. This allows one to observe, using Corollary 30 in [33, Appendix B], that a polymatroid $h$ over $N$ generates a non-modular (= tight) extreme ray of the polymatroidal cone if and only if it induces a coatom in the lattice $(\mathfrak{st}(N),\subseteq)$.
Thus, in light of Corollary 7, the computational findings from Section 5.5 imply that most of the 117 978117978117\,978117 978 extreme rays of the cone of tight polymatroids over 5-element variable set are outside the almost entropic region.
@@
which defines the almost entropic CI frame $\overline{\mathfrak{pr}}$.
{open}
Find out whether one has $\mathfrak{pr}=\overline{\mathfrak{pr}}$ or not.
It is easy to see that the frame $\overline{\mathfrak{pr}}\subseteq\mathfrak{sg}$ is closed under marginalization, copying, intersection, and lifting. Since ${\mathcal{K}}(N)\in\mathfrak{pr}(M)\subseteq\overline{\mathfrak{pr}}(M)$ if $N\subseteq M$, one, thus, has $\overline{\mathfrak{pr}}(N)\subseteq\overline{\mathfrak{pr}}(M)$ then. Therefore, in light of Definition 2.25, Task 6 is equivalent to the question whether the probabilistic CI implication on ${\mathcal{K}}(N)$ coincides with the (almost entropic) $\overline{\mathfrak{pr}}$-closure operation (for any $N$). This was basically a question raised in a private communication by our colleague Rostislav Matveev. We incline to believe that there are valid CI implications which are not true in the frame $\overline{\mathfrak{pr}}$.
The suspected probabilistic CI rules from [31, § V.B] which may appear to be examples of invalid implications in the $\overline{\mathfrak{pr}}$-frame are the rules (I:1), (I:7), (I:13), and (I:15). The reason for this guess is that to derive these implications one inevitably needs the first conditional Ingleton inequality [36] and this particular conditional inequality need not hold for almost entropic vectors [17, § V.1].
{remarks}
Let us add two remarks.
@@
    It follows from Corollary 7 that any CI implication valid in $\mathfrak{st}^{\diamond}$-frame can be viewed as a consequence of an unconstrained information inequality. Note, however, that conditional information inequalities are also studied in information theory, whose validity is restricted to and guaranteed for entropic vectors only [17]. All the so-far known probabilistic CI implications were derived from such conditional inequalities [31].

### v3

@@ 5. Self-adhesion operator and its iterating
We would like to know whether ${\mathcal{M}}\in\mathfrak{st}^{\diamond}(N)\setminus\mathfrak{st}^{\diamond\diamond}(N)$ exists in case $|N|=4$.
6. Details on computations
This section is devoted to the computational aspects of our paper: here we explain how to use SAT solvers to compute self-adhesions of suitable CI frames.
6.1. Summary from the theoretical part
We first remind the reader which of our theoretical results from previous sections are used to make the computations possible. We certainly utilized the algorithm to compute the canonical basis (Definition 2.14) described in Appendix A, which works for arbitrary Moore family ${{\mathcal{F}}}$ of subsets of a finite set $X$. Our algorithms are, however, tailored to compute the frame-based closure $\mathfrak{f}({\mathcal{M}})$ of a CI-model ${\mathcal{M}}$, where $\mathfrak{f}$ is an appropriate abstract CI frame, as introduced in Section 2.3.
In connection with our central concept of self-adhesivity, we proposed an algorithm to compute the self-adhesive closure $\mathfrak{f}^{\diamond}({\mathcal{M}})$ of a CI model ${\mathcal{M}}$ relative to the self-adhesive frame $\mathfrak{f}^{\diamond}$ (Lemma 5.1), which needs the implementation of the basic $\mathfrak{f}$-closure. The algorithm is described in Appendix B and its consistency follows from Lemma 5.4. This global algorithm utilizes as a subroutine the procedure from Lemma 4.5 to compute (conditioned) self-adhesive closure $\mathfrak{f}^{\diamond}({\mathcal{M}}|L)$ of ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ at a given subset of variables $L\subseteq N$.
6.2. The SAT problem
The Boolean satisfiability problem (SAT) deals with Boolean formulas. A Boolean formula $\phi$ is an expression using Boolean variables $V_{i}$, $i\in X$, indexed by a finite set $X$, and the Boolean operations of negation $\neg$, conjunction $\land$, and disjunction $\lor$. A mapping $a:X\to\{\texttt{true},\texttt{false}\}$ is called an assignment and it allows one to evaluate a given formula $\phi$ by replacing each variable $V_{i}$ by the Boolean value $a(i)$ and computing negations, conjunctions, and disjunctions, resulting in a Boolean value denoted by $\phi(a)$. If $\phi(a)$ is true, then $a$ is called a satisfying assignment. Denote the set of all satisfying assignments of $\phi$ by $\mathsf{SAT}(\phi)$.
Two Boolean formulas $\phi$ and $\phi^{\prime}$ are equivalent if $\mathsf{SAT}(\phi)=\mathsf{SAT}(\phi^{\prime})$. A fundamental result on Boolean formulas states that, for every formula $\phi$, there is an equivalent formula $\phi^{\prime}$ in conjunctive normal form (CNF). In this normal form, $\phi^{\prime}$ is written as a conjunction of subformulas, called clauses. Each clause, in turn, is a disjunction of literals; and each literal is either a variable or a negated variable.
The SAT problem asks to decide whether a Boolean formula given in CNF has a satisfying assignment. In this form, the problem is famously NP-complete; see [1, Section 2.3]. Variants of this problem include counting all satisfying assignments (referred to as #SAT) and enumerating all satisfying assignments (AllSAT). Remarkably efficient software for SAT, #SAT and AllSAT exists and is constantly improving thanks to regular international competitions indexed at https://www.satcompetition.org/.
6.3. Using SAT solvers for self-adhesion computation
The SAT-based modelling is a powerful tool in combinatorics. Let $X$ be the index set for variables in a Boolean formula. An assignment $a$ may be identified with the subset $\{\,i\in X:\ a(i)=\texttt{true}\,\}\subseteq X$. The set of all satisfying assignments is then an element of ${\mathcal{P}}({\mathcal{P}}(X))$. Choosing $X:={\mathcal{K}}(N)$, assignments are CI models over $N$ and the set of satisfying assignments potentially forms a CI family.
By recognizing that a CI implication $\bigwedge_{i=1}^{k}p_{i}\Rightarrow q$ (where $p_{1},\dots,p_{k},q\in{\mathcal{K}}(N)$) is nothing but a CNF clause $q\vee\neg p_{1}\vee\dots\vee\neg p_{k}$, it becomes apparent that the definition of a semi-graphoid in Section 3.1 for fixed $N$ is, in fact, a Boolean formula $\phi_{N}$ in CNF with $\mathsf{SAT}(\phi_{N})=\mathfrak{sg}(N)$. The frame of structural semi-graphoids is not defined by means of inference rules but, for each finite $N$, the Guigues–Duquenne procedure from Appendix A may be used to derive such a defining Boolean formula in conjunctive normal form.
Fix a CI family $\mathfrak{f}(N)$ over $N$ which is closed under intersection and assume that $\phi$ is a Boolean formula in CNF such that $\mathsf{SAT}(\phi)=\mathfrak{f}(N)$. We now show how the closure operator of $\mathfrak{f}(N)$ can be computed using SAT solvers. Let ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ be any set of CI statements and we fix $ab|C\not\in{\mathcal{M}}$. To test whether $ab|C\in\mathfrak{f}({\mathcal{M}})$, it suffices to test whether the CNF formula
(2) $\phi\wedge\bigwedge_{ij|K\in{\mathcal{M}}}V_{ij|K}\wedge\neg V_{ab|C}$
is satisfiable. To see this, note that the satisfying assignments are precisely those elements of $\mathfrak{f}(N)$ which contain ${\mathcal{M}}$ and not $ab|C$. If one such model exists, then $\mathfrak{f}({\mathcal{M}})$, being the intersection of all models containing ${\mathcal{M}}$, does not contain $ab|C$; otherwise the closure must contain $ab|C$ as all supersets of ${\mathcal{M}}$ in $\mathfrak{f}(N)$ contain it.
[…]
It is evident from the above description that the closure operator of $\mathfrak{f}$ is at the core of all our computations related to the self-adhesion operator. We employ a state-of-the-art incremental SAT solver to make this computation as fast as possible; all other code is less critical and mainly organizes the high-level computation, determining which closures are needed and combining the results.
6.4. Using LP solvers for self-adhesion computation
Self-adhesivity of polymatroids (see later Definition 8.1) has been used to derive further information inequalities (cf. Sections 4.1 and 8). This process follows the same pattern as our algorithms, except that Boolean formulas are replaced by polyhedral cones (whose elements are polymatroids) and, consequently, SAT solvers must be replaced by linear programming (LP) solvers. Such computations have been carried out at a large scale in [6]. The accompanying webpage
https://mathrepo.mis.mpg.de/SelfadhesiveGaussianCI/
goes into more details about the time and memory demands of SAT versus LP solvers. In the computation of self-adhesions of orientable gaussoids (using a SAT solver) versus self-adhesions of structural gaussoids (using an LP solver), the SAT solver was on average 5 times faster even though it had to deal with formulas with twice as many variables as the LP solver.
However, this comparison is unfair as the SAT solver was integrated into the process and designed to work incrementally, while the LP solver was an external program which had to be called repeatedly. About two thirds of the time it takes to compute the self-adhesive closure using the LP solver was reportedly spent on writing the linear program and starting the external LP solver. Compared to the abundance of excellent SAT solvers, it is more difficult to find a fast LP solver which is freely available software, easy to integrate into domain-specific large-scale computations and supports exact rational arithmetic which is necessary to guarantee the robustness of our results. Our computations also did not exploit recent advances in pre-processing of linear programs [19].
7. Computational results for small numbers of variables
[…]
@@ 7.2. Structural semi-graphoids over 4 variables
Note that our results agree with the analysis presented in [39, § V.A], or with the rules A3–A7 from [36]. In particular, the lattice $(\mathfrak{st}(N),\subseteq)$ is implicatively perfect in case $|N|=4$. Moreover, both the semi-graphoidal rules and the rules (E:1)–(E:5) can be generalized to any higher cardinality than $|N|\geq 4$ [3, 7].
@@ 7.3. Self-adhesive semi-graphoids over 4 variables
We also computed the canonical implicational basis ${\mathbb{G}}_{{\mathcal{F}}}$ for ${\mathcal{F}}=\mathfrak{sg}^{\diamond}(N)$ in case $|N|=4$. It decomposes into 9 permutational types: besides two semi-graphoidal types mentioned in Section 7.1 it has 7 additional types and we divided them into two subgroups. The implications in the first subgroup can be interpreted as weaker versions of probabilistic equivalences from [39, § V.A], recalled in Section 7.2, and, therefore, we use the label SE to denote them:
@@
The implications in the second subgroup can be interpreted as weaker versions of probabilistic implications presented in [39, § V.B], for which reason we use the label SI to denote them:
@@
This is a comparison with the implications from [39, § V.B]:
@@
We also confirmed independently the validity of the rules (SE:1)–(SE:2) and (SI:1)–(SI:5) within the frame $\mathfrak{sg}^{\diamond}$ by finding their certificates in the form of semi-graphoidal derivation sequences, as illustrated in Example 7.1. Specifically, we found out that the antecedent set in (SE:1) is not self-adhesive at sets $ij$ and $jk$, while the antecedent sets in (SE:2) and in the rules labelled by SI are not self-adhesive at the sets $ij$, $ik$, and $jk$.
@@ 7.4. Self-adhesive structural semi-graphoids: 4 variables
Put $N:=\{a,b,c,d\}$. The model :=${\mathcal{M}}:=\{\,ab|d,ac|d,ad|b,bc|\emptyset,bc|d\,\}$ belongs to $\mathfrak{st}^{\diamond}(N)$ while its dual model not: :=${\mathcal{M}}^{\star}:=\{\,ab|c,ac|b,ad|c,bc|a,bc|ad\,\}\not\in\mathfrak{st}^{\diamond}(N)$. The latter fact follows from the rule (SI:2) above where $i:=a$, $j:=c$, $k:=b$, and $\ell:=d$.
@@
Open problem 3.
We would like to know whether the relation $\mathfrak{sg}^{\diamond}(N)\cap\mathfrak{st}(N)=\mathfrak{st}^{\diamond}(N)$ holds in general.
7.5. Graphoids, compositional semi-graphoids, and compositional graphoids: 4 variables
A result of our computations was the finding that, in case $|N|=4$, the family $\mathfrak{gr}(N)$ contains 6482 elements, which fall down into 421 permutational types. The number of (meet) irreducible models in $\mathfrak{gr}(N)$ is 408 and they decompose into 32 types. The lattice $(\mathfrak{gr}(N),\subseteq)$ has 14 coatoms of 4 types; thus, it is not coatomistic. The family $\mathfrak{gr}^{\diamond}(N)$ in case $|N|=4$ coincides with $\mathfrak{gr}(N)$. Note that we also found that, in case $|N|=4$, precisely $369$ permutational types of graphoids are probabilistically representable.
@@ 7.6. The case of 5 variables
In general, the coatoms of this lattice correspond to the extreme rays of the cone of standardized supermodular functions, that is, to supermodular functions $m:{\mathcal{P}}(N)\to{\mathbb{R}}$ satisfying $m(K)=0$ if $K\subseteq N$, $|K|\leq 1$, see [34, Corollary 30]. Thus, the coatoms for $|N|=5$ were obtained by computing the extreme rays of that cone [33]. We raised a couple of questions:
@@
Surprising answers were that only 154 of those types belong to $\mathfrak{sg}^{\diamond}(N)$ and these coincide with the types belonging to $\mathfrak{st}^{\diamond}(N)$. This observation seems to support the conjecture indicated in Open problem 3. Consequences of those results are discussed in detail in Section 8.
@@ 8. Application to entropic region delimitation
Well-known facts are that the entropy function $h_{\mbox{\scriptsize\boldmath$\xi$}}:{\mathcal{P}}(N)\to{\mathbb{R}}$ is submodular, non-decreasing with respect to inclusion, and satisfies $h_{\mbox{\scriptsize\boldmath$\xi$}}(\emptyset)=0$. Thus, in terms of Section 3.2, it is (the rank function of) a polymatroid over $N$.
A polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ is then called entropic if $h=h_{\mbox{\scriptsize\boldmath$\xi$}}$ for some discrete random vector $\xi$ over $N$. It is called almost entropic if it is a limit of entropic polymatroids (in the Euclidean topology on ${\mathbb{R}}^{{\mathcal{P}}(N)}$). The entropic region, also named the entropy region, is the set of all entropic polymatroids and the almost entropic region is the set of all almost entropic polymatroids; the latter set is known to be a closed cone in ${\mathbb{R}}^{{\mathcal{P}}(N)}$ [41, § 15.1].
@@
Given a bijection $\varphi:N\to M$, the $\varphi$-copy of a polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ is the polymatroid $h\circ\varphi_{-1}\in{\mathbb{R}}^{{\mathcal{P}}(M)}$. A pair of sets $N,M$ is called modular in a polymatroid $\widehat{h}\in{\mathbb{R}}^{{\mathcal{P}}(NM)}$ if $\widehat{h}(N)+\widehat{h}(M)=\widehat{h}(NM)+\widehat{h}(N\cap M)$. A polymatroid $\widehat{h}\in{\mathbb{R}}^{{\mathcal{P}}(NM)}$ extends $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ if $h$ is the restriction of $\widehat{h}$ to ${\mathcal{P}}(N)$. We say that a polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ is self-adhesive at a set $L\subseteq N$ if,
@@
such that N,M is a modular pair in $~{}~{}\mbox{such that $N,M$ is a modular pair in $\widehat{h}$}\,.$
A polymatroid over $N$ is self-adhesive if it is self-adhesive at every $L\subseteq N$.
Recall from Section 3.2 that a polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ induces a CI model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ as follows: $ij|K\in{\mathcal{M}}$ if and only if $\Delta h(ij|K)=0$.
@@
Proof.
To show (a) is true, consider a set $L\subseteq N$ and the respective bijection $\varphi:N\to M$. Realize that the set of polymatroids $\widehat{h}\in{\mathbb{R}}^{{\mathcal{P}}(NM)}$ in which $N,M$ is a modular pair and that satisfy $\widehat{h}(A)=\widehat{h}(\varphi(A))$ for $A\subseteq N$ is a polyhedral cone. A polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ is self-adhesive at $L$ if and only if it belongs the projection of that set of vectors $\widehat{h}$ from ${\mathbb{R}}^{{\mathcal{P}}(NM)}$ to ${\mathbb{R}}^{{\mathcal{P}}(N)}$. Basic facts from polyhedral geometry are that the class of polyhedral cones is closed under projections and finite intersections. The application of these facts yields (a).
For (b) we first realize that any entropic polymatroid is self-adhesive. The arguments for this from [29, Corollary 1] are analogous to those in our proof of Lemma 4.7. Then (a) implies that set of self-adhesive polymatroids over $N$ is closed, which forces that it contains the closure of the entropic region.
As concerns (c), recall from the remark n. 1 in the end of Section 3.2 that ${\mathcal{M}}\in\mathfrak{st}(N)$ if and only if it is induced by a polymatroid over $N$. Given ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ induced by a self-adhesive polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$, one needs to observe that ${\mathcal{M}}$ is self-adhesive at (any) $L\subseteq N$ relative to $\mathfrak{st}$. Having fixed the corresponding bijection $\varphi:N\to M$, take the polymatroid $\widehat{h}\in{\mathbb{R}}^{{\mathcal{P}}(NM)}$ from Definition 8.1 and it is straightforward to verify that it induces ^$\widehat{{\mathcal{M}}}\in\mathfrak{st}(NM)$ which is an adhesive extension of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$. ∎
@@
Proof.
Realize that any facet-defining inequality for the cone of self-adhesive polymatroids is a valid information inequality. ∎
To interpret the results from Section 7.6 we recall some geometric facts. A polymatroid $h$ over $N$, $|N|\geq 1$, is called tight if $h(N)=h(N\setminus i)$ for any $i\in N$ and modular if $\Delta h(ij|K)=0$ for any $ij|K\in{\mathcal{K}}(N)$. Every polymatroid $h$ then decomposes into its modular part and its tight part $\widetilde{h}$ inducing the same CI model as $h$. Moreover, there exists a linear one-to-one mapping between tight polymatroids and standardized supermodular functions which preserves the induced CI model, see [34, formulas (32) in § 7.2]. This allows one to observe, using Corollary 30 in [34, Appendix B], that a polymatroid $h$ over $N$ generates a non-modular (= tight) extreme ray of the polymatroidal cone if and only if it induces a coatom in the lattice $(\mathfrak{st}(N),\subseteq)$.
Thus, in light of Corollary 8.3, the computational results from Section 7.6 imply that most of the 117 978117978117\,978117 978 extreme rays of the cone of tight polymatroids over a 5-element set are outside the almost entropic region.
@@
[…]
Open problem 4.
We would like to know whether one has $\mathfrak{pr}=\overline{\mathfrak{pr}}$ or not.
It is easy to see that the frame $\overline{\mathfrak{pr}}$ is closed under copying, marginalization, intersection, and lifting. Since ${\mathcal{K}}(N)\in\mathfrak{pr}(M)\subseteq\overline{\mathfrak{pr}}(M)$ if $N\subseteq M$, it is also closed under ascetic extension. Therefore, in light of Definition 2.20, Open problem 4 is equivalent to the question whether the probabilistic CI implication on ${\mathcal{K}}(N)$ coincides with the (almost entropic) $\overline{\mathfrak{pr}}$-closure operation (for any $N$). This was basically a question raised in a private communication by our colleague Rostislav Matveev. We incline to believe that there are valid CI implications which are not true in the frame $\overline{\mathfrak{pr}}$.
The suspected probabilistic CI rules from [39, § V.B] which may appear to be examples of invalid implications in the $\overline{\mathfrak{pr}}$-frame are the rules (I:1), (I:7), (I:13), and (I:15). The reason for this guess is that to derive these implications one inevitably needs the first conditional Ingleton inequality [42] and this particular conditional inequality need not hold for almost entropic vectors [22, § V.1].
Remarks. Let us add three comments.
@@
    It follows from Corollary 8.3 that any CI implication valid in $\mathfrak{st}^{\diamond}$-frame can be viewed as a consequence of an unconstrained information inequality. Note, however, that conditional information inequalities are also studied in information theory, whose validity is restricted to and guaranteed for entropic vectors only [22]. All the so-far known probabilistic CI implications were derived from such conditional inequalities [39].